Grade 11Physical Chemistry

A gaseous oxide contains 30.4% of nitrogen, one molecule of which contains one nitrogen atom. The density of the oxide relative to oxygen is

Himanshu
10 Years agoGrade 11

1 Answer

Vikas TU
10 Years ago
Let oxide be with one nitrogen atom given as:
NO
molar mass therfore => (16x + 14)
it is given,
14/(16x + 14)*100 =30.4
after solving,
x = 2
 
therfore the compound is: NO2
 
Relative density = 46/32

or = 23/16